The velocity of a particle which moves along



the s-axis is given by



V=2 - 4t + 58312



, Where t is in seconds and v is



in meters per second. Evaluate the position s,



velocity v, and acceleration a when t = 3 s. The



particle is at the position S, = 3 m when t=0.

1
Expert's answer
2022-03-20T18:50:59-0400

"v=2-4t+5t^2,"

"s=2t-2t^2+\frac 53t^3+c,"

"s(0)=c=3,"

"s=2t-2t^2+\frac 53 t^3+3,"

"s(3)=36~m,"

"v(3)=35~\frac m{s},"

"a=10t-4,"

"a(3)=26~\frac m{s^2}."